OpenStudy (selena2345):

Consider the pair of linear equations below: 4x+6y=12 2x+3y=6 Part A: What is the relationship, if any, between the two equations ? Part B: Does the system of equations have one solution,no solution,or infinitely many solutions? Explain. Part C: How can you verify your answers to Parts A and B by solving algebraically?

OpenStudy (anonymous):

Lines would be coincident so infinitely many solutions .....

OpenStudy (nikato):

Hint: multiply both sides of the second equation by 2. What do you get?

OpenStudy (selena2345):

can you show me ? @nikato

OpenStudy (nikato):

So multiplying 2 on the left side Gives you 2(2x+3y) Distribute and get 4x+6y Now on the right 6(2) And you get 12 So the new transformed second equation becomes 4x+6y=12 Does this make sense?

OpenStudy (anonymous):

I remember taking this lol.

OpenStudy (selena2345):

@nikato

OpenStudy (selena2345):

i need help in part a. @nikato

OpenStudy (nikato):

Ok. Now compare the transformed second equation with the original first equation. What do you see? And you didn't answer me 30 mins ago

OpenStudy (selena2345):

like terms @nikato

OpenStudy (the_fizicx99):

Look at equation A; { 4x+6y=12 } -- now look at equation B; { 2x+3y=6 } Do you see something similar?

OpenStudy (selena2345):

like terms

OpenStudy (the_fizicx99):

♦o♦ >.> No try again

OpenStudy (anonymous):

The equations have the same solutions.

OpenStudy (the_fizicx99):

X_X well that killed my moment ;-; Yes, equation A is the same as equation B because equation A is twice of equation B (2x+3y=6 )* 2 = (2x * 2) + (3y * 2) = (6*2) = 4x + 6y = 12 The same as A, they lie on top of each other >.<

OpenStudy (anonymous):

\[ 4x+6y=12 \]Can you solve for \(y\).

OpenStudy (anonymous):

@Selena2345

OpenStudy (the_fizicx99):

Or there's that method ^ :o